Simplify the complex fraction:
5
-
y
-------
2
---
y+2
Thanks,
Pat
: Simplify the complex fraction:
5
-
y
-------
2
---
y+2
Thanks,
Pat

Answer by ankor@dixie-net.com(4683)

Simplify the complex fraction:
:
5/y
-----
2/((y+2))
Invert the dividing fraction and multiply
5/y * ((y+2))/2 = (5(y+2))/(2y) = ((5y + 10))/(2y)